A 38.63-mg sample of potassium ozonide, KO3, was heated to 70 °C for 1 h, undergoing a weight loss of 7.10 mg. Write a balanced chemical reaction describing this decomposition reaction. A 29.6-mg sample of impure KO3 experiences a 4.86-mg weight loss when treated under similar condition. What is the %w/w KO3 in the sample?


1
Expert's answer
2022-02-08T11:20:03-0500

KO3----heat----> KO3+ lost weight


%w/w= (4.86/29.6)× 100


%w/w= 16.42%